OpenStudy (anonymous):

The length of a minute hand is 6". Find the length of the arc the tip sweeps going from 12-1:20 PM

OpenStudy (anonymous):

So basically every 5 minute interval is 30 degrees, so 1:20 to 12 = 360+120 degrees= 480 degrees So then s = r * theta With theta in raidans So 480 degrees to radians is 8.37758 Finally s (arc length) = r * theta = 6* 8.37758 = 50.265 Inches arc length